find the vaalue of xif p+1=0 and x^2+px-6=0
Answer by lwsshak3(11628)

find the value of x if p+1=0 and x^2+px-6=0
..
x^2+px-6=0
p+1=0
p=-1
x^2-x-6=0
(x-3)(x+2)=0
x=3
or
x=-2